Please help me I having trouble with this problem.
Business and finance.
What x number of units are sold,the price of each untit(in dollars)is given by p=-x/2+75.Find the unit price when the following quantities are sold:2,7,9,11.

Answer by malakumar_kos@yahoo.com(315)

What x number of units are sold,the price of each untit(in dollars)is given by p=-x/2+75.Find the unit price when the following quantities are sold:2,7,9,11.

given p = -x/2+75
When x = 2 then p = -2/2+75 = -1+75 = 74
When x = 7 then p = -7/2+75 = -3.5+75 = 71.5
When x = 9 then p = -9/2+75 = -4.5+75 = 70.5
When x = 11 then p = -11/2+75 = -5.5+75 = 69.5